# Hollow toroidal rotation profiles in strongly electron heated H-mode plasmas in the ASDEX Upgrade tokamak

### 05 Mar 2026 v1

@@ 4 Formation of Hollow Rotation Profiles
For these discharges, it was not possible to perform a full torque perturbation analysis, possibly because the modulation beam power was reduced too strongly ($0.7$ MW, $55$ keV) to allow sufficient penetration into the plasma core and to provide analyzable modulation amplitude profiles. Nevertheless, the time-averaged steady-state profiles provide valuable additional insight into the conditions under which hollow rotation profiles emerge.
@@
Panel (a) of Fig. 11 shows the electron density profiles for the high- and low-$n_{e}$ scenarios. The high-$n_{e}$ case closely matches the density profile of the NBI+ECRH phase of #29216 in the core within error bars (see black dashed line), while the low-fueling case exhibits an overall downward shift of the core density and modified pedestal gradients. Despite these differences, the core density gradients are found to be similar, as discussed later. Panel (b) presents the measured toroidal plasma rotation profiles. The high-$n_{e}$ case mimics the hollow rotation profile observed in #29216. In contrast, the low-$n_{e}$ case exhibits comparable edge rotation values but increases toward significantly higher rotation levels in the core with only a small hollow ”dip” around $\rho_{\varphi}\approx 0.4$. As gradients appear to be very similar between the high- and low-$n_{e}$ case, the lower overall density must lead to changes in the relative size of the transport components, preventing the formation of hollow rotation profiles. Panels (c) and (d) show the ion and electron temperature profiles, respectively. While the gradients remain very similar, a slight upward shift is observed in the low-$n_{e}$ case, consistent with the idea of identical heating power distributed over fewer particles. A comparison of the temperature profiles with those of the reference discharge shows agreement within the error bars, although some gradients appear to differ slightly. A more precise reproduction of discharge #29216 was not achieved. However, such exact agreement is not required to capture and understand the underlying dynamics associated with the density increase observed in these experiments, as discussed in the following.
@@
Next, Fig. 12 confirms the previously stated observation that the gradients are very similar between the two discharge phases, with the exception of localized ”dips” in the logarithmic ion temperature profiles, see Panel (b). These ”dips” are likely caused by disturbances in the CXRS spectra collected by the corresponding channel, combined with the limited radial smoothing applied for most accurate data analysis. Given the overall agreement with respect to experimental uncertainties in the estimation of gradients, both discharges are expected to exhibit very similar Prandtl and pinch numbers, as well as comparable intrinsic torque levels. A comparison with the NBI+ECRH phase of #29216 further shows general agreement, suggesting that similar momentum transport behavior can be expected in these new discharges.
@@
As similar dimensionless parameters are observed for all three discharges, it can be assumed that comparable transport dynamics are present in both the high- and low-$n_{e}$ reproductions of #29216. This observation supports a scenario in which a strong inward convective momentum flux, such as that inferred for the NBI+ECRH phase of #29216, see Fig. 10(b), increases significantly with the background rotation level. In the low-$n_{e}$ case, the higher pedestal-top rotation amplifies the inward convective flux, enabling it to balance the counter-current intrinsic torque. Consequently, the formation of hollow rotation profiles is largely suppressed, in agreement with the experimentally observed non-hollow rotation during the low-$n_{e}$ NBI+ECRH phase.
@@
This hypothesis can be tested by applying the transport coefficients fitted to the NBI+ECRH phase of #29216, as shown in Fig. 7, to the corresponding discharge phases of #42339 and #42340. As shown in the upper Panels of Fig. 14, and with respect to the involved error bars, this approach yields good agreement between the modeled and experimentally measured rotation profiles for both discharges over the radial range in which the original transport coefficients were fitted. The good agreement confirms that the dimensionless quantities are sufficiently similar to yield comparable transport dynamics across all three discharges. A closer analysis of the resulting momentum fluxes is shown in the lower panels. The intrinsic fluxes (solid brown line) are similar, in the high-$n_{e}$ case, the intrinsic flux responds to the $\chi_{i}$ spike at the innermost radii (see Fig. 13a), although this effect is rather localized. The same spike propagates into the diffusive flux (dotted green line), balancing the intrinsic component. Beyond this, the most significant difference is found in the convective momentum flux (dashed purple line), which increases over the entire radial range in the low-$n_{e}$ case and becomes the dominant contribution at larger radii. This suggests that the pedestal-top boundary condition, in combination with inward convective momentum transport, helps to prevent the formation of hollow rotation profiles, even in the presence of similar counter-current intrinsic torque. The black markers for both cases are very close to zero, confirming the self-consistency of the imposed combination of transport coefficients.
@@ 5 Summary
These results provide important constraints for predicting rotation behavior in future low-torque tokamak scenarios. From the perspective of future reactor operation, peaked rotation profiles with strong rotational shear are generally favorable. In the absence of strong external torque sources, this suggests that avoiding an ITG–TEM mixed turbulence regime, which can generate strong counter-current intrinsic torque, may be beneficial. However, this is practically impossible to control, since the electron and ion heat fluxes in a future reactor are comparable in magnitude, automatically leading to a mixed-mode turbulence regime. Alternatively, in an ECRH dominated reactor scenario, strong coupling between electrons and ions would allow the ion heat flux to increase relative to the electron heat flux, pushing the turbulence back toward an ITG-dominated regime.
Given the limited control over the turbulence regime in a reactor environment, strong inward convective momentum transport appears advantageous, as it can mitigate or prevent the formation of hollow rotation profiles. As demonstrated in this analysis, the effectiveness of inward convection depends critically on the overall rotation level set at the pedestal top. This motivates further investigation of mechanisms capable of generating or breaking edge intrinsic torque, such as neoclassical toroidal viscosity driven by externally applied 3D magnetic perturbations.

### 01 Apr 2026 v2

@@ 4 Formation of Hollow Rotation Profiles
For these discharges, it was not possible to perform a full torque perturbation analysis, possibly because the modulation beam power was reduced too strongly ($0.7$ MW, $55$ kV) and higher modulation frequencies were chosen, both leading to insufficient penetration into the plasma core and insufficient modulation amplitude profiles for analysis. Nevertheless, the time-averaged steady-state profiles provide valuable additional insight into the conditions under which hollow rotation profiles emerge. Future work could improve the modulation scenario, which could reduce the size of the uncertainties encountered in the analysis of the reference discharge, in particular when employing higher modulation frequencies.
@@
Panel (a) of Fig. 11 shows the electron density profiles for the high- and low-$n_{e}$ scenarios. The high-$n_{e}$ case closely matches the density profile of the NBI+ECRH phase of #29216 in the core within error bars (see black dashed line), while the low-fueling case exhibits an overall downward shift of the core density and modified pedestal gradients. Despite these differences, the core density gradients are found to be similar, as discussed later. Panel (b) presents the measured toroidal plasma rotation profiles. The high-$n_{e}$ case mimics the hollow rotation profile observed in #29216. In contrast, the low-$n_{e}$ case exhibits comparable edge rotation values but increases toward significantly higher rotation levels in the core with only a small hollow ”dip” around $\rho_{\varphi}\approx 0.4$. This can be related to the fact that the same NBI torque density acts locally on fewer particles, which are then accelerated to higher rotation values. The precise rotation level, however, is set self-consistently by the interplay of all transport mechanisms and source terms.
As gradients appear to be very similar between the high- and low-$n_{e}$ case, the lower overall density must lead to changes in the relative size of the transport components, preventing the formation of hollow rotation profiles. Panels (c) and (d) show the ion and electron temperature profiles, respectively. While the gradients remain very similar, a slight upward shift is observed in the low-$n_{e}$ case, consistent with the idea of identical heating power distributed over fewer particles. A comparison of the temperature profiles with those of the reference discharge shows agreement within the error bars, although some gradients appear to differ slightly. A more precise reproduction of discharge #29216 was not achieved. However, such exact agreement is not required to capture and understand the underlying dynamics associated with the density increase observed in these experiments, as discussed in the following.
@@
Next, Fig. 12 confirms the previously stated observation that the gradients are very similar between the two discharge phases, with the exception of localized ”dips” in the logarithmic ion temperature profiles, see Panel (b). These ”dips” are likely caused by disturbances in the CXRS spectra collected by the corresponding channel, combined with the limited radial smoothing applied for most accurate data analysis. Not shown are $q$, $\beta_{e}$, $\rho_{*}$, and $\nu_{*}$, which show moderate responses to the changed density but are otherwise very similar to those shown in Fig. 4. Given the overall agreement within experimental uncertainties in the estimation of gradients, both discharges are expected to exhibit very similar turbulence characteristics. This motivates the assumption that they could indeed exhibit very similar Prandtl and pinch numbers, as well as comparable intrinsic torque levels. A comparison with the NBI+ECRH phase of #29216 further shows general agreement, suggesting that similar momentum transport behavior can be expected in these new discharges.
@@
As similar dimensionless parameters are observed for all three discharges, see discussion of Fig. 12,it can be assumed that comparable transport dynamics are present in both the high- and low-$n_{e}$ reproductions of #29216. This observation supports a scenario in which a strong inward convective momentum flux, such as that inferred for the NBI+ECRH phase of #29216, see Fig. 10(b), increases significantly with the background rotation level. In the low-$n_{e}$ case, the higher pedestal-top rotation amplifies the inward convective flux, enabling it to balance the counter-current intrinsic torque. Consequently, the formation of hollow rotation profiles is largely suppressed, in agreement with the experimentally observed non-hollow rotation during the low-$n_{e}$ NBI+ECRH phase.
[…]
@@
This hypothesis can be tested by applying the transport coefficients fitted to the NBI+ECRH phase of #29216, as shown in Fig. 7, to the corresponding discharge phases of #42339 and #42340. As shown in the upper Panels of Fig. 14, and with respect to the involved error bars, this approach yields good agreement between the modeled and experimentally measured rotation profiles for both discharges over the radial range in which the original transport coefficients were fitted. The good agreement confirms that the dimensionless quantities are sufficiently similar to yield comparable transport dynamics across all three discharges. However, it should be noted that this does not imply a universal validity of the applied transport coefficients, but is rather a result of similar background profiles implying similar turbulence and transport characteristics.
A closer analysis of the resulting momentum fluxes is shown in the lower panels. The intrinsic fluxes (solid brown line) are similar, in the high-$n_{e}$ case, the intrinsic flux responds to the $\chi_{i}$ spike at the innermost radii (see Fig. 13a), although this effect is rather localized. The same spike propagates into the diffusive flux (dotted green line), balancing the intrinsic component. Beyond this, the most significant difference is found in the convective momentum flux (dashed purple line), which increases over the entire radial range in the low-$n_{e}$ case and becomes the dominant contribution at larger radii. This suggests that the pedestal-top boundary condition, in combination with inward convective momentum transport, helps to prevent the formation of hollow rotation profiles, even in the presence of similar counter-current intrinsic torque. The black markers for both cases are very close to zero, confirming the self-consistency of the imposed combination of transport coefficients.
@@ 5 Summary
These results provide important constraints for predicting rotation behavior in future low-torque tokamak scenarios. From the perspective of future reactor operation, peaked rotation profiles with strong rotational shear are generally favorable. In the absence of strong external torque sources, this suggests that avoiding an ITG–TEM mixed turbulence regime, which can generate strong counter-current intrinsic torque, may be beneficial. However, this is practically impossible to control, since the electron and ion heat fluxes in a future reactor are comparable in magnitude, automatically leading to a mixed-mode turbulence regime. Alternatively, in an ECRH dominated reactor scenario, strong coupling between electrons and ions would allow the ion heat flux to increase relative to the electron heat flux, pushing the turbulence back toward an ITG-dominated regime. However, the immediate implications of these insights, such as the mechanisms of intrinsic torque generation from profile-shearing effects, for integrated modeling of future devices are somewhat limited, due to the significant uncertainties associated with accurately predicting second-order profile gradients for future machines.
Therefore, given the limited control over the turbulence regime in a reactor environment, strong inward convective momentum transport appears advantageous, as it can mitigate or prevent the formation of hollow rotation profiles. As demonstrated in this analysis, the effectiveness of inward convection depends critically on the overall rotation level set at the pedestal top. This motivates further investigation of mechanisms capable of generating or breaking edge intrinsic torque. For understanding rotation levels in future low-torque devices, it is therefore important to understand edge intrinsic torque generation, such as neoclassical toroidal viscosity driven by externally applied 3D magnetic perturbations [Zhu_2006, Garofalo_2008, COLE_2011, Shaing_NF_2015], mechanisms of preferential ion losses [Stoltzfus-Dueck_PRL_2012], or the effect of strong $E\times B$ shearing [Staebler_PRL_2013, Casson_PoP_2009].